The vapour pressure of ethanol and methanol are 44.5 mm and 88.7 Hg respectively. An ideal solution is formed at the same temperature by mixing 60 g of ethanol with 40 g of methanol. Calculate the total vapour pressure of the solution and the mole fraction of methanol in the vapour.

Grade:upto college level

3 Answers

Deepak Patra
10 years ago
Hello Student,
Please find the answer to your question
Ptotal = pA + pB
Molecular weight of CH3OH = 12 + 3 + 16 + 1 = 32
Molecular weight of C­2H5OH = 24 + 5 + 16 + 1 = 46
According to Raoult’s law
Ptotal = p1 + p2
Where Ptotal = Total vapour pressure of the solution
P1 = Partial vapour pressure of one component
P2 = Partial vapour pressure of other componenet
Again, p1 = Vapour pressure (po1) * mole fraction
Similarly, p2 = Vapour pressure (po2) * mole fraction
Mole fraction of CH3OH = 40/32 /40/32+60/46 = 0.49
Mole fraction of ethanol = 60/46 /60/46+40/32 = 0.51
NOTE THIS STEP : Thus now let us first calculate the partial vapour pressures, i.e., p1 and p2 of the two component.
Partial vapour pressure of CH3OH(p1)
= 44.5 * 0.51 = 22.69 mm
∴ Total vapour pressure of the solution
= 43.48 + 22.69 mm = 66.17 mm
Mole fraction of CH3OH in vapour = 43.48/66.17 = 0.65

Thanks
Deepak patra
askIITians Faculty
Sandeep
6 years ago
P=P1+P2MOLECULAR WEIGHT OF CH3OH=32(p1)MOLECULAR WEIGHT OF C2H5OH=46(p2)NOW WE HAVEP1=vapour pressure(P`1)×mole fractionP2=vapour pressure (P`2)×mole fraction ThenMole fraction of CH3OH=0.49Mole fraction of C2H5OH=0.51P1=44.5×0.49=22.69P2=88.7×0.51=44.8P=P1+P2=66.9mm=total vapour pressure of the mixture 66.9mm Mole fraction of CH3OH in vapour is 44.8/66.17 = 0.65
Abhishat Tiwari
5 years ago
The vapor pressure of pure water is 760 mm at 25.c the vapour pressure of solution ccontanning 1(m) solution. Of glucose will be